The following information was available for Metlock, Inc. at December 31, 2017: beginning inventory $79000; ending inventory $134000; cost of goods sold $608000; and sales $888000. Metlock days in inventory in 2017 was (Round intermediate calculation to 1 decimal place, e.g. 1.2.)
Answer #1
Average inventory 106500 =(79000+134000)/2
Inventory turnover = Cost of goods sold/Average inventory
Inventory turnover 5.7 =608000/106500
Metlock days in inventory in 2017 = 365/ Inventory Turnover
Metlock days in inventory in 2017 64.0 days =365/5.7